Problem 2.

According to the study conducted last year, the average monthly consumption of senior high school students for mobile phone loads is P400. A group of statistics students believes that the amount was increased in the last quarter. Is there a reason to believe that the amount has really increased if the sample of 20 students has an average monthly expenses of P450 for mobile phone loads. Use 0.05 level of significance. Assume that the population standard deviation is P70.


1
Expert's answer
2021-06-30T03:03:21-0400

"\mu=400"

"\sigma=70"

"\bar X=450"

"n=20"

"\alpha=0.05"

"H0:\mu=400"

"Ha:\mu>400"

"z=\frac{\bar X-\mu}{\frac{\sigma}{\sqrt{n}}}"

"z=\frac{450-400}{\frac{70}{\sqrt{20}}}=3.19"

"Cv=Z_{0.95}=1.645"

3.19>1.645, thus, we reject the null hypothesis. There is enough evidence to support the claim that the average monthly consumption of senior high school has increased.